OpenStudy (anonymous):

p(w)=(3w+5)/(6-2w) do i need to solve this? or find the intercepts?

OpenStudy (jdoe0001):

what's the context?

OpenStudy (anonymous):

1. Sketch an graph of the function: . Make sure to include the asymptotes and the intercepts, along with labels and an appropriate scale. This graph must be hand drawn.

OpenStudy (jdoe0001):

ok, well, I assume you covered asymptotes already so

OpenStudy (anonymous):

I found x=3, for vertical asymptote, and y=3w+5/2(3-w) for horizontal

OpenStudy (jdoe0001):

your vertical asymptotes will occur at the roots of the denominator, when set to 0, that is 6-2w = 0 solve for "w" :) and that's your vertical asymptote

OpenStudy (jdoe0001):

6-2w =0 6 = 2w 3 = w thus yes, your vertical asymptote is at 3 :)

OpenStudy (jdoe0001):

your horizontal asymptotes occur when the numerator degree is LESS or equal to that of the denominator's if it's LESS then the horizontal asymptote occurs at the x-axis, that is y=0 if both numerator's and denominator's degree are the same, then your vertical asymptote occurs at \(\bf \cfrac{\text{leading coefficient}}{\text{leading coefficient}}\)

OpenStudy (anonymous):

yay! the rest though.. would the x intercept be at 3? and y?

OpenStudy (anonymous):

so y=7?

OpenStudy (jdoe0001):

as usual, when the function touches the x-axis, "y" is 0 when the function touches the y-axis, "x" is 0 so set either to 0 to find the other's intercept

OpenStudy (anonymous):

so is p(w) the same as y?

OpenStudy (jdoe0001):

hmm, yes

OpenStudy (jdoe0001):

what your function is using is a variable "w" as opposed to say "x" and p(w) as opposed to "y" but the axis system is generic and can use whatever you want to label the axis, so long they correlate

OpenStudy (jdoe0001):

you can if you want think of p(w)=(3w+5)/(6-2w) as y=(3x+5)/(6-2x)

OpenStudy (anonymous):

yes, i get always confused when they switch the variables-why cant life be simple?- :) thank you so much for your help, i feel i know how to, but my brain is fighting to see clearly

OpenStudy (anonymous):

so do i substitute the value of x 930 on y=(3x+5)/(6-2x) and solve, or is this not asked of me?

OpenStudy (jdoe0001):

doesn't seem so, just to get those values to graph it

OpenStudy (jdoe0001):

$$\bf p(w) = \cfrac{3w+5}{6-2w}\\ \text{y-intercept, x = 0}\\ p(w) = \cfrac{5}{6}\\ \text{x-intercept, p(w) = 0}\\ 0 = \cfrac{3w+5}{6-2w} \implies 0 = 3w+5 \implies -\cfrac{5}{3} $$

OpenStudy (jdoe0001):

and the asymptotes are at vertical, w = 3 horizontal, y= -3/2

OpenStudy (anonymous):

oh... im confused by the -5/3.. ok, thanks so much, im graphing it!

OpenStudy (jdoe0001):